### v1

# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

Abstract
We investigate the long-term relaxation of a distribution of $N$ point vortices in two-dimensional hydrodynamics, in the limit of weak collective amplification. Placing ourselves within the limit of an average axisymmetric distribution, we stress the connections with generic long-range interacting systems, whose relaxation is described within angle-action coordinates. In particular, we emphasise the existence of two regimes of relaxation, depending on whether the system’s profile of mean angular velocity (frequency) is a non-monotonic [resp. monotonic] function of radius, which we refer to as profile (1) [resp. profile (2)]. For profile (1), relaxation occurs through two-body non-local resonant couplings, i.e. $1/N$ effects, as described by the inhomogeneous Landau equation. For profile (2), the impossibility of such two-body resonances submits the system to a “kinetic blocking”. Relaxation is then driven by three-body couplings, i.e. ${1/N^{2}}$ effects, whose associated kinetic equation has only recently been derived. For both regimes, we compare extensively the kinetic predictions with large ensemble of direct $N$-body simulations. In particular, for profile (1), we explore numerically an effect akin to “resonance broadening” close to the extremum of the angular velocity profile. Quantitative description of such subtle nonlinear effects will be the topic of future investigations.
I Introduction
Two-dimensional (${2D}$) geophysical and astrophysical flows have the striking property of organizing spontaneously into coherent structures (large-scale vortices) Chavanis (2002); Bouchet and Venaille (2012). A famous example of this self-organization is Jupiter’s great red spot. Since these coherent structures result from a turbulent mixing, it has sometimes been said that “order emerges from chaos”. To account for this self-organization, Onsager Onsager (1949) considered the statistical mechanics of a system of point vortices of circulations $\gamma_{i}$ (${i\!=\!1,...,N}$) on the plane. He showed the existence of negative temperature states [^1] This was a few years before the discovery of negative temperature states in nuclear spin systems Purcell and Pound (1951). In his visionary paper, Onsager also predicted, in a famous footnote, the quantization of the circulation of point vortices in units of ${h/m}$ in superfluid helium. at which point vortices of the same sign tend to group themselves into “supervortices”. His qualitative arguments were later developed by Joyce and Montgomery Joyce and Montgomery (1973); Montgomery and Joyce (1974), Kida Kida (1975) and Pointin and Lundgren Pointin and Lundgren (1976); Lundgren and Pointin (1977) in a mean field approximation valid for ${N\!\to\!+\infty}$, with ${\gamma\!\sim\!1/N}$. Overall, this leads to the Boltzmann distribution coupled to the Poisson equation. [^2] In unpublished notes Eyink and Sreenivasan (2006), Onsager had previously developed this mean field theory by analogy with his research on electrolytes, which are also described by the Boltzmann–Poisson equations. This equilibrium state (most probable state) can be obtained by maximizing the Boltzmann entropy at fixed circulation, energy, and angular momentum. Later, Miller Miller (1990), as well as Robert and Sommeria Robert and Sommeria (1991), considered the statistical mechanics of more realistic continuous vorticity fields (instead of singular point vortices) described by the ${2D}$ Euler–Poisson equations. In particular, they derived a generalization of the Boltzmann distribution involving a continuum of vorticity levels. For a long time, the distinction between these two statistical theories was not fully understood and appreciated. It was clarified only later by developing the analogy between ${2D}$ vortices and stellar systems Chavanis et al. (1996); Chavanis (2002).
Indeed, self-gravitating systems also self-organize into coherent structures like star clusters, galaxies, or even clusters of galaxies Binney and Tremaine (2008). Stellar systems, which are fundamentally discrete (i.e. made of a collection of $N$ stars in gravitational interaction), undergo two successive types of evolution. In a first regime, one can ignore gravitational encounters between stars and can make a mean field approximation. In this collisionless regime, the locally averaged distribution function of stellar systems is governed by the Vlasov–Poisson equations. Collisionless stellar systems generically experience a process of violent relaxation towards a quasistationary (virialised) state which is a stable stationary solution of the Vlasov–Poisson equations. Lynden-Bell Lynden-Bell (1967) tried to predict this quasistationary state in terms of statistical mechanics but his prediction is hampered by the problem of incomplete relaxation Lynden-Bell (1967); Chavanis (2006). Then, on longer timescales, stellar systems evolve under the effect of gravitational encounters (finite-$N$ effects, granularities, correlations…). During this collisional regime, stellar systems have the tendency to become isothermal but their relaxation towards a true statistical equilibrium state described by the Boltzmann–Poisson equation [^3] Actually, there is no statistical equilibrium state for self-gravitating systems in ${3D}$ because the Boltzmann distribution coupled to the Poisson equation has an infinite mass. Said differently, there is no maximum entropy state at fixed mass and energy. is spoiled by physical effects such as the evaporation of high energy stars and the gravothermal catastrophe (core collapse) Lynden-Bell and Wood (1968).
Similar results apply to ${2D}$ vortices. A gas of point vortices, like a gas of stars, undergoes two successive types of evolution. In a first regime, one can ignore finite-$N$ effects and can make a mean field approximation. In this collisionless limit, the locally averaged vorticity of ${2D}$ point vortices is governed by the ${2D}$ Euler–Poisson equations. Collisionless point vortices generically experience a process of violent relaxation towards a quasistationary state which is a stable stationary solution of the ${2D}$ Euler–Poisson equations. The Miller–Robert–Sommeria Miller (1990); Robert and Sommeria (1991) statistical theory in fluid mechanics, which applies to this regime, is the counterpart of the Lynden-Bell theory Lynden-Bell (1967) in stellar dynamics. Its predictions are also affected by the problem of incomplete relaxation. Then, on longer timescales (collisional regime), ${2D}$ point vortices, just like stellar systems, evolve under the effect of “collisions” towards the statistical equilibrium state described by the Boltzmann–Poisson equation Joyce and Montgomery (1973); Montgomery and Joyce (1974); Kida (1975); Pointin and Lundgren (1976); Lundgren and Pointin (1977).
One can then develop the kinetic theory of ${2D}$ point vortices by analogy with the kinetic theory of stellar systems pioneered by Chandrasekhar Chandrasekhar (1942, 1943a, 1943b, 1949). By making a thermal bath approximation and using the methods of Brownian theory Chandrasekhar (1943c), Chandrasekhar showed that a test star experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a friction term. Similarly, under the same assumptions, a test vortex experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a drift term Chavanis (1998, 2001). As such, the drift of a point vortex is the counterpart of the dynamical friction undergone by a star. These Fokker–Planck equations relax both towards the Boltzmann distribution. In addition, the diffusion and friction (or drift) coefficients are related to each other and to the temperature through some Einstein relation. In particular, in the case of point vortices, the temperature may be negative.
However, this thermal bath approach is oversimplified and does not correctly describe the collisional evolution of the system as a whole. In particular, the Fokker–Planck equation does not conserve the total energy. A better kinetic equation for systems with long-range interactions is the Landau equation Landau (1936) or the Balescu–Lenard equation Balescu (1960); Lenard (1960) which were originally introduced for homogeneous plasmas and applied in a slightly different form to stellar systems by making a local approximation Rosenbluth et al. (1957). These kinetic equations take into account two-body correlations (collisions), so they are valid at order ${1/N}$. In the homogeneous case, they relax towards the Maxwell-Boltzmann distribution on a timescale ${(N/\ln N)\,T_{\mathrm{dyn}}}$, due to logarithmic corrections. These kinetic equations were later extended to describe the axisymmetric evolution of a point vortex gas as a whole Dubin and O’Neil (1988); Dubin (2003); Chavanis (2001); Chavanis and Lemou (2007); Chavanis (2008, 2010, 2012a, 2012b); Fouvry et al. (2016); Chavanis (2023). It turns out that their structure is seemingly different from the original (homogeneous) Landau and Balescu–Lenard equations. Indeed, they involve a resonance condition on the profile of angular velocity ${\Omega(r,t)}$. As long as the profile of angular velocity is non-monotonic, the resonance condition can be satisfied and the system relaxes through two-body effects. However, the relaxation stops when the profile of angular velocity becomes monotonic, even if the system has not reached the statistical equilibrium state. This is an example of “kinetic blocking” Chavanis (2001); Chavanis and Lemou (2007). A similar situation of kinetic blocking had been previously observed for ${1D}$ homogeneous plasmas Eldridge and Feix (1962). In that case, the Balescu–Lenard collision term vanishes identically and the evolution of the system is governed by a collision term of order ${1/N^{2}}$ that accounts for three-body correlations. An explicit expression for this collision term has been obtained recently in the case of homogeneous ${1D}$ systems with long-range interactions, when collective effects are neglected Fouvry et al. (2019, 2020). It always drives the relaxation towards the Boltzmann distribution, showing that the relaxation time scales as ${N^{2}T_{\mathrm{dyn}}}$ in that case.
In reality, stellar systems are spatially inhomogeneous. Their kinetic theory is described by the inhomogeneous Balescu–Lenard equation written with angle-action variables Heyvaerts (2010); Chavanis (2012c). When collective effects are neglected (i.e. for sufficiently hot systems), it reduces to the inhomogeneous Landau equation Chavanis (2013). These equations involve a resonance condition that depends on the orbital frequencies, ${\mathbf{\Omega}(\mathbf{J},t)}$, which themselves are functions of the actions, $\mathbf{J}$. This is similar to the kinetic equation for ${2D}$ point vortices with more general variables. For ${1D}$ inhomogeneous long-range interacting systems with a monotonic frequency profile in which only ${1\!:\!1}$ resonances are permitted, there can also be a kinetic blocking, just like for axisymmetric point vortices. In that case, one has to develop the kinetic theory at the order ${1/N^{2}}$. The expression of the collision term in the absence of collective effects has been obtained in Fouvry (2022). Except for very special interaction potentials Fouvry and Roule (2023), this equation drives the system towards the Boltzmann distribution on a timescale ${N^{2}\,T_{\mathrm{dyn}}}$.
Building upon this rich history of parallel theoretical developments, in the present work, we leverage the generality of the kinetic theory of (inhomogeneous) long-range interacting systems to apply it to describe the long-term relaxation of axisymmetric distribution of point vortices.
First, we focus on a case with a non-monotonic distribution of orbital frequency (i.e. the vortices’ angular velocity) [profile (1)]. In that case, in the limit where collective effects can be neglected, relaxation can be described with the inhomogeneous Landau equation. We compare quantitatively its prediction with measurements in direct $N$-body simulations, focusing in particular on the dependence of the relaxation rate with the total number of particles and their active fraction (i.e. the fraction of self-consistency in the mean potential). Importantly, we unveil how, close to the extremum of the angular velocity profile, the effective relaxation in the $N$-body simulations is reduced compared to the kinetic prediction. We attribute this effect to “resonance broadening” through nonlinear effects and present a heuristic correction to the Landau equation that closely matches the numerical measurements. Surprisingly, we show that as one delays the impact of nonlinearities, e.g., by increasing the number of particles and/or lowering the active fraction, the resonance broadening seems to saturate and the convergence toward the Landau prediction halts.
Second, we focus on the case of an axisymmetric distribution of point vortices with a monotonic distribution of orbital frequency. In that case, both the ${1/N}$ Landau and ${1/N}$ Balescu–Lenard equations predict a vanishing flux: a “kinetic blocking” (see, e.g., Chavanis, 2001; Chavanis and Lemou, 2007). Placing ourselves once again in the limit of weak collective amplification, we compare quantitatively measurements of the relaxation rate in numerical simulations with the prediction from the recently derived ${1/N^{2}}$ Landau equation Fouvry (2022). We obtain a satisfying agreement between both, recovering in particular the appropriate dependence with respect to the total number of particles and their active fraction.
The paper is organised as follows. In Sec. II, we detail our system. In Sec. III, we investigate long-term dynamics as driven by ${1/N}$ effects [profile (1)], while in Sec. IV, focusing on systems undergoing a kinetic blocking [profile (2)], we consider ${1/N^{2}}$ effects. We conclude in Sec. V. Throughout the main text, technical details are kept to a minimum and deferred to Appendices or to relevant references.
II System
II.1 Hamiltonian
We consider a ${2D}$ system composed of $N$ point vortices of individual circulation ${\gamma_{i}\!=\!\Gamma_{\mathrm{tot}}/N}$, with $\Gamma_{\mathrm{tot}}$ the system’s total active circulation. The dynamics of these vortices is given by the Kirchhoff–Hamilton equations (Newton, 2001)
$\gamma_{i}\,\frac{\mathrm{d}x_{i}}{\mathrm{d}t}=\frac{\partial H_{\mathrm{tot}}}{\partial y_{i}},\qquad\gamma_{i}\,\frac{\mathrm{d}y_{i}}{\mathrm{d}t}=-\frac{\partial H_{\mathrm{tot}}}{\partial x_{i}},$ (1)
with ${(x_{i},y_{i})\!=\!\mathbf{r}_{i}}$ the location of the point vortex $i$. Here, as visible in Eq. (1), $x_{i}$ and $y_{i}$ should be interpreted as canonically conjugate. The system’s total Hamiltonian is
$H_{\mathrm{tot}}=\sum_{i<j}\gamma_{i}\gamma_{j}\,U(\mathbf{r}_{i},\mathbf{r}_{j}),$ (2)
with the pairwise interaction
$U(\mathbf{r},\mathbf{r}^{\prime})=-\frac{G}{2\pi}\,\ln|\mathbf{r}\!-\!\mathbf{r}^{\prime}|$ (3)
of typical amplitude $G$. We point out that $H_{\mathrm{tot}}$ does not contain any kinetic energy term in the usual sense. We also recall that the $2D$ Newtonian potential $U$ from Eq. (3) satisfies the Poisson equation ${\Delta_{\mathbf{r}}U(\mathbf{r},\mathbf{r}^{\prime})\!=\!-G\,\delta_{\mathrm{D}}(\mathbf{r}\!-\!\mathbf{r}^{\prime})}$.
In practice, to cure possible divergences on small scales in Eq. (3), we replace the pairwise interaction with its softened analog
$U_{\epsilon}(\mathbf{r},\mathbf{r}^{\prime})=-\frac{G}{2\pi}\ln\big{(}\big{[}|\mathbf{r}\!-\!\mathbf{r}^{\prime}|^{2}\!+\!\epsilon^{2}\big{]}^{1/2}\big{)},$ (4)
with $\epsilon$ some (small) softening length. This greatly eases the implementation of the $N$-body simulations (see Appendix G).
II.2 Axisymmetric limit
When averaged over independent realisations, the system’s statistics is described via its distribution function (DF), ${F\!=\!F(\mathbf{r},t)}$, normalised to ${\!\int\!\mathrm{d}\mathbf{r}F\!=\!\Gamma_{\mathrm{tot}}}$. We assume that the mean DF, ${F({\bf r},t)\!=\!\langle\sum_{i}\gamma_{i}\delta_{\mathrm{D}}[\mathbf{r}\!-\!\mathbf{r}_{i}(t)]\rangle}$ is axisymmetric so that one has ${F\!=\!F(r,t)}$ with ${r\!=\!|\mathbf{r}|}$. Subsequently, the mean individual Hamiltonian ${H(\mathbf{r},t)\!=\!\int\mathrm{d}\mathbf{r}^{\prime}U(\mathbf{r},\mathbf{r}^{\prime})F(\mathbf{r}^{\prime},t)}$ is ${H\!=\!H(r,t)}$. We can then introduce appropriate angle-action coordinates ${(\theta,J)}$, with $\theta$ the ${2\pi}$-periodic angle and $J$ the associated action, via [^4] These follow from the generating function ${S(y,\theta)\!=\!\tfrac{1}{2}y^{2}\tan\theta}$, to be used in eq. (D.94) of Binney and Tremaine (2008).
$\theta=\tan^{-1}(x/y),\qquad J=\tfrac{1}{2}r^{2}.$ (5)
Importantly, volume conservation ensures that ${\mathrm{d}\mathbf{r}\!=\!\mathrm{d}\theta\mathrm{d}J}$. With such a choice, it is now clear how an axisymmetric distribution of point vortices can simply be interpreted as one realisation of a long-range interacting inhomogeneous systems (see, e.g., Chavanis, 2012c), characterised by some mean DF, ${F\!=\!F(J,t)}$. Along the same line, from the mean Hamiltonian, ${H\!=\!H(J,t)}$, we define the mean orbital frequency, ${\Omega(J,t)\!=\!\partial H/\partial J}$ [^5] In the original point vortex model (see, e.g., Chavanis, 2012a), the DF ${F(\mathbf{r},t)}$ corresponds to the vorticity ${\omega(\mathbf{r},t)}$, the individual Hamiltonian ${H(\mathbf{r},t)}$ corresponds to the stream function ${\psi(\mathbf{r},t)}$, and the orbital frequency ${\Omega(J,t)}$ corresponds to the angular velocity ${\Omega(r,t)}$ (for axisymmetric flows). For the $2D$ Newtonian interaction potential, one has ${\omega\!=\!-\Delta\psi}$ and ${\Omega(r,t)\!=\!\frac{1}{r^{2}}\int_{0}^{r}\mathrm{d}r^{\prime}\,r^{\prime}\omega(r^{\prime},t)}$. In Fig. 1, we illustrate the typical evolution in configuration space and action space of such axisymmetric distributions.
Refer to caption Refer to caption
Figure 1: Left: Illustration in configuration space, ${(x,y)}$, of the dynamics of a few point vortices following the DF from Eq. (6). Right: Same, but in action space ${(\theta,J)}$. In that space, the vortices follow (perturbed) straight-line trajectories.
The goal of kinetic theory is to describe the mean long-term evolution of the orbital distribution, i.e. to predict ${\partial F(J,t)/\partial t}$. Let us now describe the setup considered in this work.
II.3 Initial conditions
For the DF, we consider an initial condition of the form
$F(J)\!=\!A\bigg{(}\frac{J\!-\!J_{0}}{\sigma_{0}}\!+\!1\bigg{)}^{2}\bigg{(}\frac{J\!-\!J_{0}}{\sigma_{0}}\!-\!1\bigg{)}^{2}\!\Theta\big{[}|J\!-\!J_{0}|\!\leq\!\sigma_{0}\big{]},$ (6)
with ${\Theta[x\!\geq\!0]}$ the usual Heaviside function and ${A\!=\!15\Gamma_{\mathrm{tot}}/(32\pi\sigma_{0})}$ to ensure ${\!\int\!\mathrm{d}J\mathrm{d}\theta F\!=\!\Gamma_{\mathrm{tot}}}$. As illustrated in Fig. 2, this DF is characterised by two parameters, namely $J_{0}$ its central location and $\sigma_{0}$ its width.
Figure 2: Illustration of the DF from Eq. (6) (full lines) and the considered frequency profiles (dashed lines): Profile (1) (top panel) is non-monotonic, while Profile (2) (bottom panel) is monotonic.
This DF is non-zero only within the domain ${|J\!-\!J_{0}|\!\leq\!\sigma_{0}}$ and is continuous, along with its first derivative, on the entire action domain. As such, it is a convenient DF with which to explore long-term relaxation.
II.4 Active fraction
To limit ourselves to the dynamically hot limit, i.e. the limit of small amplification of perturbations, we assume that part of the Hamiltonian is not self-consistently generated by the vortices. As such, we introduce an active fraction, ${0\!<\!q\!<\!1}$, and make the replacement
$\gamma_{i}\to q\,\gamma_{i},$ (7)
so that the total circulation ${\Gamma_{\mathrm{tot}}\!=\!\sum_{i=1}^{N}\!\gamma_{i}}$ is also reduced by a factor $q$. Following this modification, we add an external potential to our system to ensure that systems with different $N$ and $q$ share the exact same mean-field orbital frequencies.
More precisely, we replace the total Hamiltonian from Eq. (2) with
$H_{\mathrm{tot}}=\sum_{i=1}^{N}\gamma_{i}\,U_{\mathrm{ext}}(\mathbf{r}_{i})+\sum_{i<j}^{N}\gamma_{i}\,\gamma_{j}\,U_{\epsilon}(\mathbf{r}_{i},\mathbf{r}_{j}).$ (8)
In that expression, $U_{\epsilon}$ is the softened pairwise interaction from Eq. (4). We also introduced an external potential, $U_{\mathrm{ext}}$, that reads
$U_{\mathrm{ext}}=H_{0}-H_{\epsilon}[F].$ (9)
In that expression, ${H_{\epsilon}[F](\mathbf{r})\!=\!\!\int\!\mathrm{d}\mathbf{r}^{\prime}F(\mathbf{r}^{\prime},t\!=\!0)\,U_{\epsilon}(\mathbf{r},\mathbf{r}^{\prime})}$ is the (softened) potential generated by the initial DF from Eq. (6). We also introduced ${H_{0}\!=\!H_{0}(J)}$ an axisymmetric external potential. Importantly, with the present choice, at ${t\!=\!0}$, we have ${\langle H_{\mathrm{tot}}\rangle\!=\!\sum_{i}\gamma_{i}\,H_{0}(J)}$, i.e. the initial mean-field orbital frequencies are simply ${\Omega(J)\!=\!\mathrm{d}H_{0}/\mathrm{d}J}$. From it, we also define the dynamical time as ${T_{\mathrm{dyn}}\!=\!2\pi/|\Omega(J_{0})|}$. With this setup, changing the value of $N$ amounts to changing the amplitude of the Poisson fluctuations, while changing $q$ amounts to changing the dynamical temperature, i.e. the level of self-consistent amplification in the system. Importantly, changing $N$ or changing $q$ does not lead to any change in the mean frequency profile. In the following, for simplicity, we assume that the point vortices have the same circulation, ${\gamma_{i}\!=\!\gamma}$.
II.5 Frequency profiles
For the frequency profiles, we consider two different cases, profiles (1) and (2), as illustrated in Fig. 2. As detailed in Appendix A, these correspond to the (unsoftened) potential generated by a simple axisymmetric background distribution (Eq. 37). [^6] We emphasise that the orbital frequency (angular velocity) profile ${\Omega(J)}$ is not self-consistently generated by the point vortices that constitute our system, i.e. by the DF from Eq. (6), but is imposed from the outside, e.g., being generated by a background distribution of point vortices exterior to our system. This allows us to prescribe the orbital frequency in a convenient manner. The main difference between these two frequency profiles, ${J\!\mapsto\!\Omega(J)}$, is that profile (1) is non-monotonic while profile (2) is monotonic. As the upcoming section will detail, this plays a crucial role in defining the system’s regime of relaxation.
III ${1/N}$ dynamics
III.1 Kinetic equation
First, let us consider profile (1). At order ${1/N}$, relaxation is driven by two-body effects. In the limit where collective effects can be neglected, this is described by the ${1/N}$ inhomogeneous Landau equation (see, e.g., Chavanis, 2013). In the present context, it reads
$\frac{\partial F(J,t)}{\partial t}=\gamma\frac{\partial}{\partial J}\bigg{[}{}$ $\sum_{k}k\!\!\int\!\!\mathrm{d}J_{1}\,\big{|}\Lambda_{k}(\mathbf{J})\big{|}^{2}$
[…]
The Landau equation (Eq. 10) satisfies a couple of important conservation laws. Up to prefactors, these are
[…]
If we define the Boltzmann entropy as
$S(t)=-\!\!\int\!\!\mathrm{d}J\,F(J,t)\,\ln\big{[}F(J,t)\big{]},$ (12)
the ${1/N}$ Landau equation satisfies an $H$-theorem, i.e. it drives ${\mathrm{d}S/\mathrm{d}t\!\geq\!0}$. Finally, the thermodynamical equilibria, i.e. the Boltzmann DFs, of the form
$F_{\mathrm{B}}(J)\propto\mathrm{e}^{-\beta H(J)+\gamma J},$ (13)
are generic steady states of the ${1/N}$ Landau equation, i.e. they satisfy ${\partial F_{\mathrm{B}}/\partial t\!=\!0}$ when inserted into Eq. (10). There is no guarantee, however, that the inhomogeneous Landau equation at order ${1/N}$ relaxes towards the Boltzmann distribution because of (local) kinetic blockings, as explained below.
Performing explicitly the sum over $k$ in Eq. (10), the ${1/N}$ Landau equation can be rewritten as
$\frac{\partial F(J,t)}{\partial t}={}$ $\gamma\frac{\partial}{\partial J}\bigg{[}\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{\mathrm{D}}\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}$ (14)
[…]
with
$\big{|}\Lambda_{\mathrm{tot}}(\mathbf{J})\big{|}^{2}\!=\!2\sum_{k>0}k\,|\Lambda_{k}(\mathbf{J})|^{2}$ (15)
the total coupling coefficient. In that form, Eq. (14) involves the simple resonance condition ${\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]}$. If the frequency profile is monotonic, the condition ${\Omega(J_{1})\!=\!\Omega(J)}$ implies ${J_{1}\!=\!J}$ and Eq. (10) vanishes, i.e. ${\partial F/\partial t\!=\!0}$. There is no relaxation through ${1/N}$ effects. This is a situation of global kinetic blocking at order ${1/N}$. In the presence of a non-monotonic frequency profile, for a given $J$, there may exist another ${J_{1}\!\neq\!J}$ such that ${\Omega(J_{1})\!=\!\Omega(J)}$. In that case, Eq. (10) does not vanish and the system can relax through ${1/N}$ effects. This is the case for profile (1), as visible in Fig. 2.
In that case, assuming that collective effects are negligible, we find from Eq. (14) that the relaxation time, i.e. the time for $F$ to change by order unity, scales like
$T_{\mathrm{relax}}\simeq T_{\mathrm{dyn}}\,N/q^{2},$ (16)
with respect to the total number of particles, $N$, and the system’s active fraction, $q$. As a result, it is useful to introduce the rescaled relaxation rate
$\mathcal{R}_{1}(J,t)=\frac{\partial N(\!<\!J,t)/\partial t}{N/T_{\mathrm{dyn}}}\,\frac{N}{q^{2}},$ (17)
with ${N(\!<\!J,t)\!=\!(2\pi/\gamma)\int_{0}^{J}\mathrm{d}J^{\prime}F(J^{\prime},t)}$ the number of particles with an action smaller than $J$ at time $t$. Its time derivative is given by ${\partial N(\!<\!J,t)/\partial t\!=\!(2\pi/\gamma)\mathcal{F}(J,t)}$, with the Landau flux ${\partial F(J,t)=\partial\mathcal{F}(J,t)/\partial J}$, following Eq. (14). Here, $\mathcal{R}_{1}$ is a dimensionless quantity, but more importantly, following Eq. (16), the ${1/N}$ Landau equation predicts that $\mathcal{R}_{1}$ should be independent of both $N$ and $q$.
III.2 Linear stability
Let us now ensure that we consider systems that are dynamically hot enough. As detailed in Appendix C, the strength of collective effects can be estimated via the “dielectric matrix”, ${\mathbf{E}_{k}(\omega_{\mathrm{R}})}$ for ${\omega_{\mathrm{R}}\!\in\!\mathbb{R}}$. Then, the dressed coupling coefficients, ${\Lambda^{\mathrm{d}}_{k}}$, that appear in the ${1/N}$ Balescu–Lenard equation when compared with the ones from the Landau equation scale typically like ${\Lambda^{\mathrm{d}}_{k}(\mathbf{J};\omega_{\mathrm{R}})\!\sim\!\Lambda_{k}(\mathbf{J})/|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$, with ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}})]}$, the determinant of the dielectric matrix. In practice, the closer ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$ is from unity, the weaker are collective effects. On the contrary, if ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!0}$, then the system supports an infinite amplification, i.e. this is a neutral mode. Importantly, since it describes collisionless response theory, we note that ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$ is independent of $N$ and depends only on the active fraction, $q$.
In Fig. 3, we illustrate the Nyquist contour for profile (1), i.e. the curve of ${\omega_{\mathrm{R}}\!\mapsto\!|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$ in the complex plane.
Figure 3: Nyquist contour for profile (1) for various values of the active fraction $q$, for the resonance ${k\!=\!1}$. We refer to Appendix C for the definition of the “dielectric matrix”, ${\mathbf{E}_{k}(\omega)}$. For ${q\!\gtrsim\!0.0010}$, the system is found to be linearly unstable.
In that figure, we observe that for ${q\!\gtrsim\!0.0010}$, the system is linearly unstable since the Nyquist contour encloses the origin of the complex plane. [^7] See, e.g., Chavanis (2012d) for a simple illustration of the Nyquist method in homogeneous systems with long-range interactions. The possibly unstable profile (1) looks similar to the “gravity” case (compare Fig. 3 with fig. 2 of Chavanis (2012d)), while the always stable profile (2) looks similar to the “plasma” case (compare Fig. 10 with fig. 21 of Chavanis (2012d)). Moreover, we find that for ${q\!\lesssim\!0.0003}$, the contour remains close to ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!1}$, i.e. collective amplification is small. In all the upcoming explorations, we limit ourselves to numerical simulations with ${q\!\leq\!0.0003}$, i.e. well within the range of applicability of the ${1/N}$ Landau equation in which collective effects are neglected.
III.3 Small-scale contributions
Before comparing quantitatively the kinetic prediction with $N$-body simulations, let us consider the contributions from small scales in Eq. (10). Indeed, that equation involves an infinite sum over the resonance numbers $k$, so that one should ensure that the limit ${k\!\to\!+\infty}$ is meaningful.
Following Eq. (15), we can naturally decompose the relaxation rate from Eq. (17) as
$\mathcal{R}_{1}(J)\!=\!\sum_{k>0}\mathcal{R}_{1}(J,k),$ (18)
where we dropped the dependence with respect to $t$ to shorten the notations. In Fig. 4, we illustrate the function ${k\!\mapsto\!\mathcal{R}_{1}(J\!=\!1.1\,J_{0},k)}$, emphasising that the considered action satisfies a non-local resonance condition (Fig. 2).
Figure 4: Contribution to the relaxation rate, ${\mathcal{R}_{1}(k)}$ (Eq. 18), evaluated in ${J/J_{0}\!=\!1.1}$, as a function of the resonance number, $k$. The different thicknesses correspond to different values of the softening length, $\epsilon$. For small $k$, we recover the expected scaling in $k^{-1}$.
From Fig. 4, we can make three observations: (i) For $k$ small enough, one finds ${\mathcal{R}_{1}(k)\!\propto\!k^{-1}}$, in agreement with the decay rate that can directly be inferred from Eq. (14), as detailed in Appendix D. (ii) As $k$ increases, we find that ${k\!\mapsto\!\mathcal{R}_{1}(k)}$ decreases faster than $k^{-1}$. This comes from the fact that for $k$ large enough, the difference between the two actions that are in resonance leads to a decay faster than algebraic (see, e.g., Eq. 45b). (iii) Finally, as one increases the softening length, $\epsilon$, (Eq. 4), the decay rate with respect to $k$ gets even faster. This makes perfect sense physically, since softening naturally washes out contributions from small scales. Overall, in the present regime, the ${1/N}$ Landau equation does not suffer from any small-scale divergence, contrary to the small-scale Coulomb logarithm that plagues the ${1/N}$ Landau equation when applied to ${3D}$ self-gravitating systems (see, e.g. Chavanis, 2013).
III.4 Relaxation rate
We are now set to compare the prediction from Eq. (10) with measurements in $N$-body simulations. This is presented in Fig. 5.
Figure 5: Initial relaxation rate in action space, $\mathcal{R}_{1}$, for profile (1). For the simulations, the uncertainty is illustrated with the 16% and 84% contours obtained by bootstrap over the available realisations (see Appendix G). For the prediction, the full line corresponds to the ${1/N}$ Landau equation (Eq. 10), while the dashed line is from the “regularised” ${1/N}$ Landau equation (Eq. 25), with the regularised time, ${T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!307}$, adjusted to the slope around the sign switch of $\mathcal{R}_{1}$. Close to the extremum of the frequency profile, broadening effects play a crucial role.
A couple of remarks can be made on Fig. 5. For now, let us focus on the ${1/N}$ kinetic prediction given by Eq. (10). First, for ${J/J_{0}\!\lesssim\!0.9}$, it predicts ${\mathcal{R}_{1}\!=\!0}$. Glancing back at Fig. 2, this corresponds to a range of actions, $J$, for which the resonance condition, ${\Omega(J_{1})\!=\!\Omega(J)}$ (Eq. 14), imposes ${J_{1}\!=\!J}$. As such, only local resonances are permitted by the mean frequency profile, hence leading to the local vanishing of Eq. (14) because of the antisymmetry of the collision kernel. Therefore, at order ${1/N}$, this region does not relax towards the Boltzmann distribution, because of a local kinetic blocking. As such, to relax towards the statistical equilibrium state, $1/N^{2}$ effects must be taken into account, as studied in the next section for a purely monotonic frequency profile. Then, for ${J/J_{0}\!\simeq\!1.03}$, we find that the kinetic prediction for $\mathcal{R}_{1}$ abruptly changes sign. This is due to the extremum of the frequency profile that appears at this location, as visible in Fig. 2. Although the kinetic prediction is discontinuous near this extremum, it suffers from no divergences.
We can now compare this prediction with the measurement from the $N$-body simulations. Reassuringly, we recover that for ${J/J_{0}\!\lesssim\!0.9}$, relaxation is greatly delayed, a local kinetic blocking. Second, as one considers larger action, say ${0.9\!\lesssim\!J/J_{0}\!\lesssim\!1.0}$, we find a good agreement between the numerical simulations and the kinetic prediction from the ${1/N}$ Landau equation. However, as ${J}$ nears ${1.03\,J_{0}}$, i.e. as $J$ nears the extremum of the frequency profile, we find that the kinetic prediction overestimates the numerical measurements. In particular, the simulations do not exhibit the discontinuity predicted by the Landau equation, and present a rather smooth change of sign. This mismatch is to be attributed to the limitations of some of the hypotheses made in the derivation of the ${1/N}$ Landau equation. In Fig. 5, we present a prediction with a “regularised” Landau equation, that offers a much more satisfying agreement with the numerical simulation. This is the topic of the next section.
III.5 Resonance broadening
The ${1/N}$ Landau equation (Eq. 14) involves the sharp resonance condition, ${\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]}$. Mathematically speaking, this sharp resonant Dirac comes from the fact that the derivation of the Landau equation requires the propagation of the dynamics of the DF’s fluctuations, at linear order, up to infinite dynamical times (see, e.g., Chavanis, 2012c). Yet, as a result of non-linearities and diffusion, such an assumption is only partially true. This effect can be responsible in particular for “resonance broadening” (see, e.g. Weinstock, 1969), as we will now briefly explore, in a heuristic fashion.
[…]
$\frac{\partial\delta F}{\partial t}+[\delta F,H]+[F,\delta H]=0,$ (19)
with $H$ the system’s mean Hamiltonian, and ${\delta H\!=\!\delta H[\delta F]}$ the fluctuations in the Hamiltonian generated by ${\delta F}$. When considered within angle-action coordinates, Eq. (19) is straightforward to solve and gives a solution of the form
$\delta F_{k}(J,t)\propto\delta F_{k}(J,0)\,\mathrm{e}^{-\mathrm{i}k\Omega(J)t},$ (20)
with ${\delta F_{k}}$ the Fourier transform in angle of ${\delta F}$, for the harmonic $k$. Physically, this solution assumes that the DF fluctuations follow the system’s unperturbed mean field orbits. Finally, by sending the dynamical time to infinity, we find that
$\mathrm{Re}\bigg{[}\lim\limits_{T\to+\infty}\!\!\int_{0}^{T}\!\!\mathrm{d}t\,\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{R}}t}\bigg{]}\simeq\pi\,\delta_{\mathrm{D}}(\omega_{\mathrm{R}}),$ (21)
with ${\omega_{\mathrm{R}}\!=\!\Omega(J)\!-\!\Omega(J_{1})}$ the resonance frequency. Therefore, in Eq. (21), we recover, heuristically, the sharp resonance condition from Eq. (14). To sum up, the ${1/N}$ Landau equation involves a Dirac delta as a result of two assumptions (i) the DF fluctuations are propagated linearly, i.e. along the unperturbed orbits; (ii) and this holds for infinitely long dynamical times.
Yet, these two assumptions cannot hold for so long, since we neglected the contributions from the nonlinear term, ${[\delta F,\delta H]}$, in Eq. (19). Heuristically, let us therefore assume that Eq. (20) must be weighted down by an exponential damping factor on some “regularisation” timescale, $T_{\mathrm{reg}}$. The dynamics of the DF fluctuations would become
$\delta F_{k}(J,t)\!\propto\!\delta F_{k}(J,0)\,\mathrm{e}^{-\mathrm{i}k\Omega(J)t}\,\mathrm{e}^{-kt/T_{\mathrm{reg}}}.$ (22)
Physically, this means that after a time ${T_{\mathrm{reg}}/k}$, fluctuations on scale $k$ are scattered away from their mean field orbit. This prevents the further build up of resonant interactions. Performing the same integral as in Eq. (21), we obtain
$\mathrm{Re}\bigg{[}\lim\limits_{T\to+\infty}\!\!\int_{0}^{T}\!\!\mathrm{d}t\,\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{R}}t}\,\mathrm{e}^{-t/T_{\mathrm{reg}}}\bigg{]}\simeq\pi\,\delta_{T_{\mathrm{reg}}}(\omega_{\mathrm{R}}),$ (23)
where we introduced the Lorentzian
$\delta_{T_{\mathrm{reg}}}(\omega)=\frac{1}{\pi}\frac{T_{\mathrm{reg}}}{1+(\omega T_{\mathrm{reg}})^{2}}.$ (24)
Naturally, in the limit ${T_{\mathrm{reg}}\!\to\!+\infty}$, one has ${\delta_{T_{\mathrm{reg}}}(\omega)\!\to\!\delta_{\mathrm{D}}(\omega)}$, hence recovering a sharp resonance condition.
Armed with this regularised resonance condition, we can then, heuristically, introduce “resonance broadening” in Eq. (14) by replacing its Dirac delta with this broadened condition. It becomes
$\frac{\partial F(J,t)}{\partial t}={}$ $\gamma\frac{\partial}{\partial J}\bigg{[}\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{T_{\mathrm{reg}}}\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}$ (25)
[…]
Of course, we stress that Eq. (25) is only a heuristic kinetic equation, since, for example, it does not conserve the total energy, contrary to Eq. (14).
In Fig. 6, we compute the relaxation rate predicted by Eq. (25), as one varies the regularisation timescale, $T_{\mathrm{reg}}$.
Figure 6: Predicted relaxation rate for profile (1) given by Eq. (25), as one varies the regularisation time $T_{\mathrm{reg}}$, using the same convention as in Fig. 5. For ${T_{\mathrm{reg}}\!\to\!+\infty}$, one recovers the ${1/N}$ Landau prediction from Eq. (14).
In that figure, we find that broadening effects smooth out the discontinuity predicted by the ${1/N}$ Landau equation (Eq. 14) close to the extremum of the frequency profile. Reassuringly, we also find that as ${T_{\mathrm{reg}}\!\to\!+\infty}$, the regularised prediction converges to the (sharp) Landau prediction.
Armed with a regularised kinetic equation (Eq. 25), we may use it to compare with $N$-body simulations. This is first presented in Fig. 5, where we adjusted the considered $T_{\mathrm{reg}}$ to match the slope of the numerical simulations, near the frequency extremum (Appendix D). Interestingly, we find that Eq. (25), although heuristic, offers a satisfactory match with the numerical measurements, over the whole action domain. This highlights the fact that nonlinear broadening effects are likely responsible for the effective smoothing of the relaxation rate near the extremum of the orbital frequency.
In practice, in Fig. 5, we presented a single value of ${(N,q)}$. Let us now explore the dependence of this broadening as one varies the total number of particles, $N$, and the system’s active fraction, $q$. In Fig. 7, we explore the dependence of the relaxation rate, $\mathcal{R}_{1}$, as one varies $N$ and $q$.
Figure 7: Relaxation rate for profile (1), using the same convention as in Fig. 5, as one varies $q$ for fixed $N$ (top), or varies $N$ for fixed $q$ (bottom). Here, we represent only the median value over the available realisations. As one increases $N$ and decreases $q$, the relaxation rate increases, but only up to some saturation.
First, in that figure, following the rescaling from Eq. (17), we find that the dependence with respect to ${(N,q)}$ is only very mild. This emphasises that, indeed, ${1/N}$ effects are responsible for this system’s relaxation. Yet, as one increases $N$ and/or lowers $q$, one can tentatively note some increase in the overall amplitude of the measured relaxation rate. However, this increase is only limited and ultimately saturates. As such, even if one keeps increasing $N$ and/or lowering $q$, the relaxation rate, $\mathcal{R}_{1}$, saturates and stops converging further toward the discontinuous prediction from the ${1/N}$ Landau equation (see Fig. 5).
Let us now be more specific on this (tentative) saturation effect. For a given value of ${(N,q)}$, let us therefore estimate the system’s nonlinear timescale, $T_{\mathrm{nl}}$. This time would then be the appropriate regularisation timescale, $T_{\mathrm{reg}}$, to be considered in Eq. (25) when comparing with numerical simulations. Heuristically, the nonlinear time, $T_{\mathrm{nl}}$, corresponds to the duration during which the quasilinear assumption from Eq. (19) holds. In that view, let us therefore equate the contribution from the nonlinear term ${[\delta F,\delta H]}$, with the rate of change of the DF fluctuation, ${\partial_{t}\delta F}$, in Eq. (19). We find that ${|\delta F|/T_{\mathrm{nl}}\!\simeq\!G\,|\delta F|^{2}}$, where we assumed that ${|\delta H|\!\simeq\!G\,|\delta F|}$. Assuming that the statistics of the DF fluctuations is Poissonian (see, e.g., Eq. 41 in Chavanis, 2012c), we can take ${|\delta F|^{2}\!\simeq\!\gamma\Gamma_{\mathrm{tot}}}$, with ${\gamma\!=\!\Gamma_{\mathrm{tot}}/N}$ the vortices’ individual circulations, and ${\Gamma_{\mathrm{tot}}}$ involving the system’s active fraction. Overall, we find the scaling
$T_{\mathrm{nl}}\simeq\frac{\sqrt{N}}{q}\,T_{\mathrm{dyn}}.$ (26)
Physically, the dependence of Eq. (19) makes sense. Indeed, by increasing $N$, i.e. by reducing the level of the Poisson fluctuations, and by lowering $q$, i.e. by reducing the level of collective amplification, one increases the system’s nonlinear time, i.e. one allows for the quasilinear assumption to hold longer in Eq. (19).
Interestingly, taking inspiration from Taylor and McNamara (1971); Dubin (2003), one can offer another heuristic derivation of the scaling of $T_{\mathrm{nl}}$. Rather than being driven by nonlinearities affecting the dynamics of ${\delta F}$, let us assume that it is the (finite-$N$ driven) diffusion of $F$ that drives itself the broadening of the resonance condition in Eq. (25). Dropping all factors of order unity, and keeping only the scaling dependence with respect to $N$ and $q$, one can, heuristically, rewrite Eq. (25) as
$\frac{\partial F}{\partial t}\propto\frac{q^{3}}{N}\,\frac{T_{\mathrm{nl}}}{1\!+\!(\Delta\Omega\,T_{\mathrm{nl}})^{2}},$ (27)
where the factor ${q^{3}/N}$ comes from ${\gamma F(J)F(J_{1})}$ in Eq. (25), ${\Delta\Omega\!=\!\Omega(J)\!-\!\Omega(J_{1})}$ is the resonance frequency, and we expanded the broadened Dirac, $\delta_{T_{\mathrm{reg}}}$, using Eq. (24) and imposing ${T_{\mathrm{nl}}\!=\!T_{\mathrm{reg}}}$. Let us then assume that it is the change in $F$ described by Eq. (27) that is itself responsible for resonance broadening. Hence, we may make the replacement ${\partial F/\partial t\!\to\!q/T_{\mathrm{nl}}}$ in Eq. (27), so that it becomes
$\frac{1}{T_{\mathrm{nl}}/T_{\mathrm{dyn}}}\simeq\frac{q^{2}}{N}\,\frac{T_{\mathrm{nl}}/T_{\mathrm{dyn}}}{1\!+\!(\Delta\Omega\,T_{\mathrm{nl}})^{2}}.$ (28)
Equation (28) is a self-consistent relation for the nonlinear time, $T_{\mathrm{nl}}$ (see, e.g., Taylor and McNamara, 1971; Dubin, 2003). This self-consistent relation is explored in much more detail in Appendix H. Close to the extremum of the frequency profile, we may take ${\Delta\Omega\,T_{\mathrm{nl}}\!\to\!0}$, so that Eq. (28) gives
$\frac{1}{(T_{\mathrm{nl}}/T_{\mathrm{dyn}})^{2}}\simeq\frac{q^{2}}{N}.$ (29)
Ultimately, we recover the exact same scaling as in Eq. (26). Of course, it is no surprise that both approaches lead to the same scaling, since the kinetic equation is, by construction, sourced by nonlinear effects, i.e. one has ${\partial F/\partial t\!=\!-\langle[\delta F,\delta H]\rangle}$.
We can now use numerical simulations to investigate the validity of the previous heuristic interpretations, as one varies the values of ${(N,q)}$. For a given ${(N,q)}$, we can first compute the system’s nonlinear time, $T_{\mathrm{nl}}$, following Eq. (26). For the same parameters, we can also rely on numerical simulations to measure the system’s relaxation rate, $\mathcal{R}_{1}$. Then, just like in Fig. 5, we can adjust for the value of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 25), to best reproduce the slope of the relaxation rate near the extremum of the frequency profile (following Fig. 14), i.e. where broadening effects play the most important role. In Fig. 8, we explore the dependence of $T_{\mathrm{reg}}$ as a function of $T_{\mathrm{nl}}$ for a large swath of values of ${(N,q)}$.
Figure 8: Dependence of the regularisation time, $T_{\mathrm{reg}}$ (Eq. 25), as a function of the nonlinear time, $T_{\mathrm{nl}}$ (Eq. 26), using numerical simulations with different values of $N$ (different colors) and $q$ (different symbols). Here, $T_{\mathrm{reg}}$ has been obtained following Appendix D to best reproduce the relaxation rate near the frequency extremum. Surprisingly, for large values of $T_{\mathrm{nl}}$, we find that $T_{\mathrm{reg}}$ saturates, i.e. the measured relaxation rate stops converging to the (sharp) kinetic prediction from Eq. (14).
First, in Fig. 8, for small values of $T_{\mathrm{nl}}$, we find that increasing $T_{\mathrm{nl}}$, indeed, leads to an increase in $T_{\mathrm{reg}}$. Phrased differently, in systems with strong resonance broadening, delaying the appearance of nonlinear effects leads to a more peaked relaxation rate near the frequency extremum. Unfortunately, the recovery of the expected trend ${T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}}$ is only tentative here. This is due, in particular, to the reduced range in $T_{\mathrm{nl}}$ that was considered because of the numerical complexity of simulating the present system of point vortices, for a large number of pairs ${(N,q)}$ and of independent realisations (see Appendix G for details).
Second, in Fig. 8, we find that for all systems with ${T_{\mathrm{nl}}\!\gtrsim\!3\!\times\!10^{5}\,T_{\mathrm{dyn}}}$, the inferred regularisation timescale, $T_{\mathrm{reg}}$, seems to saturate. Phrased differently, even if one keeps increasing the value of $N$ and/or decreasing the value of $q$, the measured relaxation rate stops converging further toward the (discontinuous) prediction from the ${1/N}$ Landau equation (Eq. 14). This is a very surprising result, and one for which we currently have no convincing explanation. Naturally, one could be worried that this saturation behaviour stems from some issues in the numerical simulations. To limit possible concerns, we performed a couple of convergence tests, as detailed in Appendix G. To summarize, we empirically find that
$\frac{T_{\mathrm{reg}}}{T_{\mathrm{dyn}}}\simeq\begin{cases}\displaystyle\propto\frac{\sqrt{N}}{q}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\lesssim 3\!\times\!10^{5},\\[8.61108pt] \displaystyle 3\!\times\!10^{2}{}&\displaystyle\mathrm{if}\quad\frac{\sqrt{N}}{q}\gtrsim 3\!\times\!10^{5}.\end{cases}$ (30)
Of course, the various pieces of evidence presented here only offer some glimpse on the role of nonlinear effects and resonance broadening near frequency extrema, i.e. near the breakdown of the ${1/N}$ Landau equation. At this stage, we can already list three possible venues to improve upon the results presented in the present section: (i) One should go beyond the heuristic regularisation of Eq. (25) to derive a self-consistent kinetic equation in that regime; (ii) One should reproduce the effect of resonance broadening in long-range interacting systems with much cheaper computations, e.g., with the quasiperiodic cube (see, e.g., Magorrian, 2021) or with classical Heisenberg spins (see, e.g. Barré and Gupta, 2014). These would be useful to check numerically the expected scaling ${T_{\mathrm{reg}}\!\propto\!T_{\mathrm{nl}}}$; (iii) Finally, leveraging the same systems, one should explore the relaxation of systems with much larger values of $T_{\mathrm{nl}}$ to confirm the saturation effect observed in Fig. 8. All these investigations are left for future work.
III.6 Long-term relaxation
Ultimately, the present system relaxes toward the Boltzmann distribution (Eq. 13). And, we refer to Appendix E for details on how this equilibrium may be determined given some initial conditions.
In Fig. 9, we use $N$-body simulations to illustrate this long-term relaxation to the statistical equilibrium.
Figure 9: Relaxation in $N$-body simulations of profile (1) toward the thermodynamical equilibrium (Eq. 13). Different colors correspond to different times (here the median value over the available realisations), while the dashed line is the expected Boltzmann equilibrium (see Appendix E). Relaxation is delayed by a local kinetic blocking induced by the absence of resonances in certain regions of the non-monotonic frequency profile (see Fig. 2).
In that figure, one can note that the relaxation of the system toward the central region is particularly slow. This follows from the particular shape of the present non-monotonic frequency profile (see Fig. 2). Indeed, particles within the system’s inner regions, i.e. with small actions, cannot resonantly couple to more external regions, i.e. to particles with larger actions. This was responsible for the local kinetic blocking, already visible in the ${1/N}$ Landau flux from Fig. 5. As such, these inner regions cannot relax through $1/N$ effects and must wait for the outer regions to diffuse, via ${1/N^{2}}$ effects, so that non-local resonances get populated and that the more efficient ${1/N}$ relaxation may proceed. This lack of efficient resonant couplings is responsible for the system’s particularly slow diffusion within the core regions. Phrased differently, although the system’s frequency profile is non-monotonic, there still exist regions in action space that only relax, at least initially, via ${1/N^{2}}$ effects. Such next-order contributions are the focus of the upcoming section.
IV ${1/N^{2}}$ dynamics
IV.1 Kinetic equation
We now consider profile (2), i.e. a system with a monotonic frequency profile (Fig. 2). In that case, the resonance condition in the ${1/N}$ Landau equation (Eq. 10) imposes ${J_{1}\!=\!J}$, hence leading to the vanishing of the crossed term ${\mathbf{k}\!\cdot\!\partial F_{2}/\partial\mathbf{J}}$. This is a kinetic blocking, and there is no relaxation driven by ${1/N}$ effects. [^8] Kinetic blocking at order ${1/N}$ also holds when accounting for collective effects, as described by the inhomogeneous Balescu–Lenard equation. Relaxation is greatly hampered, and occurs only through ${1/N^{2}}$ effects. Placing themselves within this regime and neglecting collective effects, Fouvry (2022) derived a closed kinetic equation describing this relaxation. The inhomogeneous ${1/N^{2}}$ Landau equation reads
$\frac{\partial F(J,t)}{\partial t}=\gamma^{2}\frac{\partial}{\partial J}\bigg{[}{}$ $\sum_{k_{1},k_{2}}\!(k_{1}\!+\!k_{2})\mathcal{P}\!\!\int\!\!\mathrm{d}J_{1}\mathrm{d}J_{2}\,\big{|}\Lambda_{k_{1}k_{2}}(\mathbf{J})\big{|}^{2}$
[…]
with $\mathcal{P}$ the Cauchy principal value acting on the integral over ${\mathrm{d}J_{1}}$. We refer to Appendix F for the expression of the coupling coefficients, ${|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}}$, which are independent of both $N$ and $q$. In Eq. (31), we shortened the notations with the 3-vectors ${\mathbf{J}\!=\!(J,J_{1},J_{2})}$, ${\mathbf{\Omega}\!=\!(\Omega[J],\Omega[J_{1}],\Omega[J_{2}])}$, and ${\mathbf{k}\!=\!(k_{1}\!+\!k_{2},-k_{1},-k_{2})}$, as well as ${F_{3}(\mathbf{J})\!=\!F(J)F(J_{1})F(J_{2})}$. At the present time, the generalisation of Eq. (31) that accounts for collective effects is unknown.
Just like the ${1/N}$ Landau equation (Eq. 10), the ${1/N^{2}}$ Landau equation (Eq. 31) conserves the total circulation, momentum and energy (Eq. 11), satisfies an $H$-theorem for the Boltzmann entropy (Eq. 12), and vanishes exactly for the Boltzmann DF (Eq. 13).
Equation (31) contains an intricate resonance condition, ${\delta_{\mathrm{D}}(\mathbf{k}\!\cdot\!\mathbf{\Omega})}$, involving three particles. Even with a monotonic frequency profile, this resonance condition allows for non-local resonances, i.e. triplets of actions, $\mathbf{J}$, that are not all identical. As a result, no further kinetic blocking is generically to be expected. [^9] Though, we refer to Fouvry and Roule (2023) for the (involved) example of a bare second-order kinetic blocking for a particular class of potentials of interaction. For the potential of interaction of point vortices (Eqs. 3 and 4), there is no kinetic blocking at the order $1/N^{2}$ and the DF relaxes through Eq. (31) towards the Boltzmann DF. In that case, we find from Eq. (31) that the relaxation time scales like
$T_{\mathrm{relax}}\simeq T_{\mathrm{dyn}}\,N^{2}/q^{4},$ (32)
with respect to $N$ and $q$, in the limit where collective effects can be neglected. As a result, it is useful to introduce the rescaled relaxation rate
$\mathcal{R}_{2}(J,t)=\frac{\partial N(\!<\!J,t)/\partial t}{N/T_{\mathrm{dyn}}}\,\frac{N^{2}}{q^{4}},$ (33)
using the same notation as in Eq. (17). In that case, the ${1/N^{2}}$ Landau equation (Eq. 31) predicts that $\mathcal{R}_{2}$ should be independent of both $N$ and $q$.
IV.2 Linear stability
The scaling from Eq. (32) only holds in the limit of weak collective amplification. Once again, we can use linear response theory to determine an appropriate range of active fraction for which collective effects are negligible. This is illustrated in Fig. 10 for profile (2) using Nyquist contours.
Figure 10: Same as Fig. 3 but for profile (2). Here, for all values of $q$, the system is always found to be linearly stable.
Interestingly, in this figure, we find that the geometry of the Nyquist contours is such that, whatever the considered value of $q$, the contour never encloses the origin of the complex plane. Phrased differently, contrary to profile (1), here for profile (2), whatever the system’s active fraction, the system’s initial distribution is linearly stable. Of course, linear stability is not synonymous with weak collective amplification. Indeed, in Fig. 10, we find that for ${q\!\gtrsim\!0.004}$, the Nyquist contour gets extremely close to the origin, i.e. there exists real frequencies for which the system amplifies very significantly perturbations. In practice, in all the upcoming simulations, we limit ourselves to simulations with ${q\!\leq\!0.0007}$, so that Eq. (31) applies.
IV.3 Small-scale contributions
As for the ${1/N}$ case, let us first focus on the contributions from small scales. Indeed, Eq. (31) involves an infinite sum over the resonance numbers ${(k_{1},k_{2})}$, i.e. one ought to ensure that the limits ${k_{1},k_{2}\!\to\!+\infty}$ are meaningful.
Following Eq. (31), we decompose the relaxation rate from Eq. (33) as
$\mathcal{R}_{2}(J)=\sum_{k>0}\mathcal{R}_{2}(J,k),$ (34)
where we introduced
$\mathcal{R}_{2}(J,k)=\sum_{\mathclap{\begin{subarray}{c}k_{1},k_{2}\\ \mathrm{Max}[|k_{1}|,|k_{2}|]=k\end{subarray}}}\mathcal{R}_{2}(J,k_{1},k_{2}).$ (35)
In Fig. 11, we illustrate the dependence of ${k\!\mapsto\!\mathcal{R}_{2}(J\!=\!1.1\,J_{0},k)}$.
Figure 11: Same as Fig. 4, but for the ${1/N^{2}}$ Landau equation (Eq. 31) and profile $(2)$ (see Eq. 34). For small $k$, i.e. scales larger than $\epsilon$, we recover the expected scaling in $k^{-2}$.
From that figure, one can make a couple of observations: (i) For $k$ small enough, one finds ${\mathcal{R}_{2}(k)\!\propto\!k^{-2}}$. This is in agreement with the decay rate that can directly be inferred from Eq. (31), as detailed in Appendix F. (ii) As one increases the softening length, $\epsilon$, the contribution to the relaxation rate decays faster with $k$. As expected, introducing softening leads to the vanishing of the contribution from small scales. Overall, we find that Eq. (31) does not suffer from any small scale divergence, when applied to the dynamics of point vortices.
IV.4 Relaxation rate
We are now set to compare the kinetic prediction from Eq. (31) with measurements in $N$-body simulations. This is presented in Fig. 12.
Figure 12: Same as Fig. 5 but for profile (2). For that profile, relaxation is driven by ${1/N^{2}}$ effects (Eq. 31), and hence is much delayed. The origin of the mismatch for large actions remains unclear.
In that figure, we find a welcome agreement between the ${1/N^{2}}$ prediction and the numerical measurements. This is one of the main result of the present work, i.e. a quantitative validation of the ${1/N^{2}}$ kinetic theory of point vortices. One should point out the presence of a mismatch for ${J/J_{0}\!\gtrsim\!1.1}$. Such a disagreement could stem from contributions associated with collective effects, since Eq. (31) neglects them altogether. We point out that all previous explicit applications of the ${1/N^{2}}$ kinetic equation have always uncovered a mismatch, see, in particular, fig. 2 in Fouvry et al. (2019) for the HMF model, fig. 1 in Fouvry et al. (2020) for the ring model, and fig. 1 in Fouvry (2022) for Heisenberg spins. A next step to clarify this effect would be to generalise Eq. (31) to the case with collective effects, i.e. to derive, for the first time, the ${1/N^{2}}$ Balescu–Lenard equation.
Building upon the previous case of profile (1), one could be worried that resonance broadening effects could be polluting the measurement from Fig. 12. In Fig. 13, we repeat the measurement of the relaxation rate, $\mathcal{R}_{2}$, for various values of $N$ and $q$.
Figure 13: Same as Fig. 13, but for profile (2). As one varies $N$ and/or $q$, the (rescaled) relaxation rate, $\mathcal{R}_{2}$ (Eq. 33), remains, in essence, unchanged.
In that figure, contrary to Fig. 7, we observe no clear trend with respect to modifications of $N$ and/or $q$. As a result, for the present monotonic frequency profile, there does not seem to be any clear signature from nonlinear effects that would cause a significant broadening of the three-body resonance condition from Eq. (31).
V Conclusions
In this paper, we investigated the long-term relaxation of long-range interacting point vortices in ${2D}$ hydrodynamics within axisymmetric configurations. We showed how the generic kinetic theory of inhomogeneous long-range interacting systems can readily be applied to point vortices. In that context, we emphasised the existence of two regimes of relaxation: (i) if the system’s mean frequency profile is non-monotonic, the distribution of point vortices relaxes through ${1/N}$ effects. In the limit of negligible collective amplification, this is governed by the ${1/N}$ inhomogeneous Landau equation (Eq. 10); (ii) if the mean frequency profile is monotonic, the system is then submitted to a kinetic blocking, which prevents ${1/N}$ effects from driving any relaxation. In that case, neglecting once again collective effects, the system’s relaxation is described by the ${1/N^{2}}$ inhomogeneous Landau equation (Eq. 31).
For both regimes of relaxation, we reported on the agreement between the prediction from kinetic theories and measurements in direct numerical simulations. Using these same simulations, we also recovered the expected scalings with respect to the total number of particles and the system’s active fraction. Finally, we emphasised the signature of nonlinear resonance broadening near the extremum of the frequency profile.
Naturally, the present work is only one small step toward building a finer understanding of long-term relaxation. Let us conclude this work by mentioning a few possible venues for future investigations.
${1/N^{2}}$ Balescu–Lenard equation. At order ${1/N}$, when collective effects cannot be neglected, the Landau equation becomes the Balescu–Lenard equation (Heyvaerts, 2010; Chavanis, 2012c). In the presence of a kinetic blocking, the ${1/N^{2}}$ version of the Balescu–Lenard equation is currently unknown. Although this is no easy challenge, it would be worthwhile to derive such a generalisation, to investigate the impact of collective effects in kinetic blockings.
Resonance broadening. Of course, it would be interesting to investigate further the effect of resonance broadening unveiled in Fig. 5. On the numerical front, one could revisit this problem using $N$-body simulations in a simpler toy model. For example, the homogeneous Hamiltonian Mean Field model Antoni and Ruffo (1995) with the kinetic energy, ${\tfrac{1}{2}v^{2}}$, changed for ${\tfrac{1}{3}v^{3}}$, hence ensuring a non-monotonic mean frequency profile, similar to the one from profile (1) here (see Fig. 2). Similarly, the case of classical Heisenberg spins could also prove insightful (see, e.g. Barré and Gupta, 2014). On the theoretical front, going beyond the heuristic regularisation from Eq. (25) is particularly challenging, in particular to ensure the expected conservation laws and $H$-theorem. One possible venue could be to explore systematic methods stemming from renormalisation theory (see, e.g., Krommes, 2002, and references therein).
Large deviations. We focused our interest on the ensemble-averaged evolution. Going beyond the mean dynamics, one should also investigate the dispersion of independent realisations around that mean evolution. This is the promise of large deviation theory applied to kinetic equations Feliachi and Bouchet (2022); Feliachi and Fouvry (2024), though its quantitative comparison with $N$-body simulations remains to be done.
Marginal stability. In the presence of a monotonic frequency profile, an axisymmetric distribution of point vortices is always submitted to a kinetic blocking, independently of the strength of collective effects. Yet, as a system nears marginal stability, i.e. as the oscillation of the damped modes take longer and longer to fade away, mode-particle couplings are expected to get more and more efficient at driving the system’s relaxation (see, e.g., Rogister and Oberman, 1968; Hamilton and Heinemann, 2020). It would be interesting to investigate the impact that such “quasilinear” contributions can have on kinetic blockings.
Spectrum of circulations. Here, we considered systems with a single population of vortices, sharing all the exact same individual circulation. In the view of better understanding segregation processes, one should investigate systems with different circulations, in particular different signs (see, e.g., Dubin, 2003; Chavanis, 2023).
Acknowledgements.
This work is partially supported by the grant Segal ANR-19-CE31-0017 of the French Agence Nationale de la Recherche and by the Idex Sorbonne Université. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We warmly thank S. Flores and A. El Rhirhayi for detailed remarks and suggestions. We are grateful to S. Rouberol for the smooth running of the Infinity cluster where the simulations were performed.
Appendix A Background profiles
In this Appendix, we detail our prescription for the external potentials introduced in Eq. (9).
The mean self-consistent potential generated by the point vortices reads
$H_{\epsilon}[F]{}$ $=\!\!\int\!\!\mathrm{d}\mathbf{r}^{\prime}\,U_{\epsilon}(\mathbf{r},\mathbf{r}^{\prime})\,F(r^{\prime})$
$=2\pi\!\!\int\!\!\mathrm{d}J^{\prime}\,U_{\epsilon,k=0}(J,J^{\prime})\,F(J^{\prime}),$ (36)
with ${F(J^{\prime})}$ the DF from Eq. (6), and the Fourier-transformed (softened) interaction potential, ${U_{k,\epsilon}(J,J^{\prime})}$ following Eq. (44). Fortunately, this integral, as well as the associated orbital frequency, can be explicitly performed, using symbolic computing (Mathematica), even in the presence of softening. We implemented these expressions in the $N$-body simulations (see Appendix G).
For profile (1), we assume that it is generated by the background DF
$F_{\mathrm{b}}(J)=\Theta[J\!\geq\!J_{\mathrm{b}}]\,\frac{\Gamma_{\mathrm{tot}}^{\mathrm{b}}}{\pi\sigma_{\mathrm{b}}}\,\frac{1}{[1+|J\!-\!J_{\mathrm{b}}|/\sigma_{\mathrm{b}}]^{3}},$ (37)
with ${\Gamma_{\mathrm{tot}}^{\mathrm{b}}\!=\!\Gamma_{\mathrm{tot}}}$ the background total circulation. Here, $J_{\mathrm{b}}$ is the shift of the Lorentzian, while $\sigma_{\mathrm{b}}$ is its width. Using once again Eq. (36), one can compute explicitly the associated, unsoftened, potential. Although somewhat lengthy, this expression can be obtained using symbolic computing. In practice, in the $N$-body simulations, we imposed ${J_{\mathrm{b}}/J_{0}\!=\!\tfrac{1}{2}}$ and ${\sigma_{\mathrm{b}}/J_{0}\!=\!1}$, for which the generated frequency profile is non-monotonic in the vicinity of the active DF, see Fig. 2.
For profile (2), we consider Eq. (37) in the limit ${J_{\mathrm{b}}\!=\!0}$. In that case, the (unsoftened) mean Hamiltonian reads
$H_{0}(J)=\frac{G\Gamma_{\mathrm{tot}}^{\mathrm{b}}}{4\pi}\,\frac{1-(1\!+\!J/\sigma_{\mathrm{b}})\,\ln[2(J\!+\!\sigma_{\mathrm{b}})]}{1\!+\!J/\sigma_{\mathrm{b}}},$ (38)
while the mean frequency is ${\Omega_{0}(J)\!=\!\Omega_{\mathrm{b}}\,g(J/\sigma_{\mathrm{b}})}$ with the frequency scale ${\Omega_{\mathrm{b}}\!=\!-G\Gamma_{\mathrm{tot}}^{\mathrm{b}}/(2\pi\sigma_{\mathrm{b}})}$ and the dimensionless monotonic function
$g(x)=\frac{1\!+\!\tfrac{1}{2}x}{(1\!+\!x)^{2}}.$ (39)
Such a model is a pleasant choice as it allows for the resonance condition in Eq. (31) to be explicitly solved. Indeed, for ${0\!<\!\varpi\!\leq\!1}$, one has
$g(x)=\varpi\;\;\Longrightarrow\;\;x=\frac{1\!-\!4\varpi\!+\!\sqrt{1\!+\!8\varpi}}{4\varpi}.$ (40)
This greatly eases the numerical implementation of the ${1/N^{2}}$ kinetic equation. In practice, in the $N$-body simulations, we imposed ${\sigma_{\mathrm{b}}/J_{0}\!=\!1}$.
Appendix B Interaction potential
In the absence of any softening, the pairwise interaction potential (Eq. 3) reads
$U(\mathbf{r},\mathbf{r}^{\prime})=-\frac{G}{2\pi}\,\ln|\mathbf{r}\!-\!\mathbf{r}^{\prime}|.$ (41)
Its Fourier expansion in angle is
$U(\mathbf{r},\mathbf{r}^{\prime})=\sum_{k=-\infty}^{+\infty}U_{k}(J,J^{\prime})\,\mathrm{e}^{\mathrm{i}k(\theta-\theta^{\prime})},$ (42)
with ${U_{k}(J,J^{\prime})\!=\!U_{k}(r[J],r^{\prime}[J^{\prime}])}$ following (see eq. (A15) in Chavanis, 2012a)
$U_{0}(r,r^{\prime}){}$ $=-\frac{G}{2\pi}\,\ln(r_{\mathrm{max}}),$ (43a)
$U_{k}(r,r^{\prime})$ $\,=\frac{G}{4\pi}\frac{1}{|k|}\,\bigg{(}\frac{r_{\mathrm{min}}}{r_{\mathrm{max}}}\bigg{)}^{|k|}.$ (43b)
In that expression, we introduced ${r_{\mathrm{min}}\!=\!\mathrm{Min}[r,r^{\prime}]}$ and ${r_{\mathrm{max}}\!=\!\mathrm{Max}[r,r^{\prime}]}$. Using the relation ${J\!=\!\tfrac{1}{2}r^{2}}$, it is then straightforward to obtain the coupling coefficient in action space, ${U_{k}(J,J^{\prime})}$.
To soften the interaction on small scales, we transform Eq. (41) into
$U(\mathbf{r},\mathbf{r}^{\prime})=-\frac{G}{2\pi}\,\ln\big{(}\big{[}|\mathbf{r}\!-\!\mathbf{r}^{\prime}|^{2}\!+\!\epsilon^{2}\big{]}^{1/2}\big{)},$ (44)
with $\epsilon$ some fixed softening length. Following the same approach as in appendix B of Weinberg (1986), the Fourier expansion from Eq. (43) becomes
$U_{0}(r,r^{\prime}){}$ $=-\frac{G}{2\pi}\,\ln(r_{\alpha}),$ (45a)
$U_{k}(r,r^{\prime}){}$ $=\frac{G}{4\pi}\frac{1}{|k|}\bigg{(}\frac{r_{\beta}}{r_{\alpha}}\bigg{)}^{|k|},$ (45b)
where we introduced the two radii
$r_{\alpha}{}$ $=\bigg{[}\tfrac{1}{2}\bigg{(}\big{\{}\big{(}(r\!+\!r^{\prime})^{2}\!+\!\epsilon^{2}\big{)}\,\big{(}(r\!-\!r^{\prime})^{2}\!+\!\epsilon^{2}\big{)}\big{\}}^{1/2}$ (46a)
$\quad\quad\quad+r^{2}\!+\!r^{\prime 2}\!+\!\epsilon^{2}\bigg{)}\bigg{]}^{1/2},$
$r_{\beta}{}$ $=\frac{rr^{\prime}}{r_{\alpha}}.$ (46b)
Appendix C Linear response theory
Balescu–Lenard equation. When accounting for collective effects, the inhomogeneous Landau equation (Eq. 10) becomes the inhomogeneous Balescu–Lenard (BL) equation (Heyvaerts, 2010; Chavanis, 2012c). It has the exact same functional form except that the bare coupling coefficients, ${U_{k}(J,J^{\prime})}$, become the (frequency-dependent) dressed ones, ${U^{\mathrm{d}}_{k}(J,J^{\prime};\omega)}$, to be used in Eq. (56).
Self-consistent relation. The dressed coupling coefficients satisfy the self-consistent relation (see, e.g., Chavanis, 2012c)
$U^{\mathrm{d}}_{k}(J,J^{\prime};{}$ $\omega)=U_{k}(J,J^{\prime})$ (47)
[…]
with ${\Omega^{\prime\prime}\!=\!\Omega(J^{\prime\prime})}$. We also introduced the regularised complex Hilbert kernel (see, e.g., Fouvry and Prunet, 2022, and references therein)
$\frac{1}{z_{-}}=\lim\limits_{T\to+\infty}\frac{1-\mathrm{e}^{-\mathrm{i}zT}}{z}.$ (48)
The dressed coupling coefficients, ${U^{\mathrm{d}}_{k}(\omega)}$, capture the strength of the collective amplification in the system.
Matrix rewriting. In inhomogeneous systems, ${U^{\mathrm{d}}_{k}(\omega)}$, is typically evaluated using Kalnajs’ basis method (Kalnajs, 1976). Here, we resort to a simpler method that frees us from having to construct tailored basis elements, as we now detail.
First, in Eq. (47), we note that both the resonance number, $k$, and the frequency, $\omega$, are external labels of the integral over ${\mathrm{d}J^{\prime\prime}}$. Let us therefore assume that we can decompose the action domain into a finite number of bins of length ${\Delta J}$. This provides us then with a list of action, ${\{J_{i}\}_{i}}$, around which each bin is centered. We now introduce the matrices, $\mathbf{U}_{k}$ and ${\mathbf{U}^{\mathrm{d}}_{k}(\omega)}$, via
$\big{(}\mathbf{U}_{k}\big{)}_{ij}{}$ $=U_{k}(J_{i},J_{j}),$ (49a)
$\big{(}\mathbf{U}^{\mathrm{d}}_{k}(\omega)\big{)}_{ij}{}$ $=U^{\mathrm{d}}_{k}(J_{i},J_{j};\omega).$ (49b)
Approximating the integral from Eq. (47) with the midpoint rule, it becomes the matrix relation
$\mathbf{U}^{\mathrm{d}}_{k}(\omega)=\mathbf{U}_{k}+\mathbf{U}_{k}\,\mathbf{M}_{k}(\omega)\,\mathbf{U}^{\mathrm{d}}_{k}(\omega),$ (50)
where we introduced the diagonal matrix
[…]
After manipulation, Eq. (50) finally becomes
$\mathbf{E}_{k}(\omega)\,\mathbf{U}^{\mathrm{d}}_{k}(\omega)=\mathbf{U}(k),$ (52)
or equivalently
$\mathbf{U}^{\mathrm{d}}_{k}(\omega)=\mathbf{E}_{k}(\omega)^{-1}\,\mathbf{U}(k),$ (53)
where we introduced the “dielectric” matrix
$\mathbf{E}_{k}(\omega)=\mathbf{I}-\mathbf{U}(k)\,\mathbf{M}_{k}(\omega),$ (54)
with $\mathbf{I}$ the identity matrix. Heuristically, Eq. (52) should be interpreted as ${U^{\mathrm{d}}\!\sim\!U/E}$, hence describing collective dressing. In particular, for a given resonance number, $k$, a system supports a mode at frequency $\omega_{\mathrm{M}}$ if one has
$\det\big{[}\mathbf{E}_{k}(\omega_{\mathrm{M}})\big{]}=0.$ (55)
Importantly, the present approach (i) does not require any potential/density basis elements; (ii) can readily be applied to softened pairwise interactions.
Nyquist contours. In Figs. 3 and 10, the Nyquist contours represent, in the complex plane, the values of the function ${\omega_{\mathrm{R}}\!\in\!\mathbb{R}\!\to\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)]\!=\!|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|}$. Phrased differently, these figures show the imaginary part of ${|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|}$ as a function of its real part, as one varies $\omega_{\mathrm{R}}$. In practice, (i) we used $500$ points to sample the action domain in Eq. (49), and (ii) we regularised Eq. (51) with the small positive imaginary part ${\eta\!=\!10^{-3}\,\Omega(J_{0})}$, so that ${1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)_{-}\!\to\!1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)}$. We checked that varying the number of sampling points or the value of $\eta$ had no impact on the plotted Nyquist curves.
Appendix D ${1/N}$ kinetic theory
${1/N}$ Landau equation. Following Chavanis (2012c), in the absence of collective effects, the inhomogeneous Landau equation (Eq. 10) satisfies
$\big{|}\Lambda_{k}(\mathbf{J})\big{|}^{2}=2\pi^{2}\,\big{|}U_{k}(J,J_{1})\big{|}^{2},$ (56)
where ${U_{k}(J,J_{1})}$ is the Fourier transformed interaction potential (see Appendix B). Performing the sum over $k$, one readily obtains Eq. (14). For the unsoftened interaction potential, the sum can be carried out explicitly (see, e.g., appendix B in Chavanis, 2010).
Small-scale contributions. In Fig. 4, we present the respective contribution of small scale resonances, i.e. large $k$, to the ${1/N}$ Landau relaxation rate. Let us recover the main features of that figure directly from Eq. (10). For ${J^{\prime}\!\to\!J}$, Eq. (45b) gives ${U_{k}(J,J^{\prime})\!\simeq\!k^{-1}}$, so that Eq. (56) gives ${|\Lambda_{k}(\mathbf{J})|^{2}\simeq k^{-2}}$. When used in Eq. (15), we find from Eq. (14) that each scale contributes like $k^{-1}$. In practice, this scaling only holds for $k$ small enough, before the difference between the two resonating actions $J$ and $J_{1}$ leads a decay faster than $k^{-1}$ in Eq. (45b). Similarly, increasing the considered softening length accelerates, as expected, the decay of the small-scale contributions. These various trends are all recovered in Fig. 4.
Computing the kinetic prediction. In practice, to evaluate Eq. (14), one invokes the relation
$\delta_{\mathrm{D}}\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}=\sum_{J_{\star}}\frac{\delta_{\mathrm{D}}(J\!-\!J_{\star})}{|\mathrm{d}\Omega/\mathrm{d}J_{\star}|},$ (57)
where the $J_{\star}$ satisfy ${\Omega(J_{\star})\!=\!\Omega(J)}$. Used in Eq. (14), we get
$\frac{\partial F(J,t)}{\partial t}=\gamma\frac{\partial}{\partial J}\bigg{[}\sum_{J_{\star}}{}$ $\frac{\big{|}\Lambda_{\mathrm{tot}}(J,J_{\star})\big{|}^{2}}{|\mathrm{d}\Omega/\mathrm{d}J_{\star}|}$ (58)
[…]
We implemented this expression to obtain Fig. 5, limiting the sum over $k$ in Eq. (15) to ${|k|\!\leq\!100}$.
Computing the regularised prediction. To compute prediction from the “regularised” ${1/N}$ Landau equation (Eq. 25), we evaluated the integral over ${\mathrm{d}J_{1}}$ therein using a simple midpoint rule with 1 000 nodes within the interval ${J_{0}\!-\!\sigma_{0}\!\leq\!J_{1}\!\leq\!J_{0}\!+\!\sigma_{0}}$, limiting, once again, the sum over $k$ in Eq. (15) to ${|k|\!\leq\!100}$.
Inferring the regularisation time. In Fig. 8, we compare the value of nonlinear time, $T_{\mathrm{nl}}$ (Eq. 26), with the regularisation time, $T_{\mathrm{reg}}$, inferred from $N$-body simulations. To obtain this figure, we proceeded as follows. First, we computed the regularised relaxation rate from Eq. (25) for ${T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!10\!\times\!2^{n}}$ with ${0\!\leq\!n\!\leq\!7}$. Then, for each such prediction, we performed a linear fit of the relaxation rate within the domain ${1.022\!\leq\!J/J_{0}\!\leq\!1.042}$, the action domain centered around the change of sign of the relaxation rate. These measurements are presented in Fig. 14.
Figure 14: Dependence of the predicted central slope of the relaxation rate with respect to the action (see, e.g., Fig. 6), as a function of the regularisation time, $T_{\mathrm{reg}}$, used in Eq. (25).
In that figure, we find a simple power-law dependence linking the regularisation time, $T_{\mathrm{reg}}$, with the central slope of the relaxation rate as a function of the action. Finally, we used the measurements from Eq. (14) to infer the “best” regularisation time presented in Fig. 8.
Appendix E Boltzmann equilibrium
The present systems have three global invariants. Following Eq. (11), these are (i) the total circulation $\Gamma$; (ii) the total momentum $L$; and (iii) the total energy $E$. In practice, since we consider systems with small active fraction, $q$, for simplicity, we may approximate the total energy by limiting ourselves only to the axisymmetric external potential, $H_{0}$, introduced in Eq. (9). Doing so, we neglect the “self-gravitating” contribution from the particles’ pairwise couplings, and the total energy (Eq. 11c) simply becomes ${E\!=\!\!\int\!\mathrm{d}J\,H_{0}(J)\,F(J,t)}$.
As a result of these conservations, the associated Boltzmann equilibria generically read (Eq. 13)
$F_{\mathrm{B}}(J)=\alpha\,\mathrm{e}^{-\beta H_{0}(J)+\gamma J},$ (59)
where ${(\alpha,\beta,\gamma)\!=\!\mathbf{X}}$ are three parameters to be found to comply with the conservation of ${(\Gamma,L,E)\!=\!\mathbf{Y}}$. Phrased differently, given some initial DF, ${F_{0}\!=\!F(t\!=\!0)}$, the associated long-term thermodynamical Boltzmann equilibrium is specified by the set of parameters, $\mathbf{X}$, such that ${\mathbf{Y}[F_{\mathrm{B}}(\mathbf{X})]\!=\!\mathbf{Y}[F_{0}]}$.
Such a self-consistent relation can be solved using Newton’s algorithm. Given some guess $\mathbf{X}^{(n)}$, we perform the iteration
$\mathbf{X}^{(n+1)}=\mathbf{X}^{(n)}-\mathsf{J}^{-1}[\mathbf{X}^{(n)}]\,\big{(}\mathbf{Y}[\mathbf{X}^{(n)}]\!-\!\mathbf{Y}_{0}\big{)},$ (60)
with ${\mathbf{Y}_{0}\!=\!\mathbf{Y}[F_{0}]}$. In that expression, we also introduced the Jacobian matrix
$\big{(}\mathsf{J}[\mathbf{X}]\big{)}_{ij}=\frac{\partial Y_{i}[\mathbf{X}]}{\partial X_{j}}.$ (61)
One can then iterate Eq. (60) up to convergence, i.e. up to ${\max_{i}|X_{i}^{(n+1)}\!-\!X_{i}^{(n)}|\!\leq\!\epsilon}$, with, here, ${\epsilon\!=\!10^{-12}}$.
To obtain the Boltzmann DF represented in Fig. 9, we computed numerically the integrals defining the invariants $\mathbf{Y}$ (Eq. 11) as well as the Jacobian $\mathsf{J}$ (Eq. 61) using the midpoint rule with $10^{5}$ nodes within the range ${0\!\leq\!J/J_{0}\!\leq\!10^{3}}$. The initial guess was fixed to ${\mathbf{X}^{(0)}\!=\![q\Gamma_{\mathrm{tot}}^{\mathrm{b}}/(2\pi),0,0]}$. Convergence was reached in 20 iterations, leading to the final value ${\mathbf{X}^{(+\infty)}\!\simeq\![8.88\!\times\!10^{-4}/\Gamma_{\mathrm{tot}}^{\mathrm{b}},-40.0/(G\Gamma_{\mathrm{tot}}^{\mathrm{b}}),6.26\!\times\!10^{-3}/J_{0}]}$.
Appendix F ${1/N^{2}}$ kinetic theory
${1/N^{2}}$ Landau equation. Following Fouvry (2022), the coupling coefficients of the inhomogeneous ${1/N^{2}}$ Landau equation read
$\big{|}\Lambda_{k_{1}k_{2}}(\mathbf{J})\big{|}^{2}\!=\!2\pi^{3}\bigg{|}\frac{\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}\mathcal{U}^{(1)}_{k_{1}k_{2}}\!(\mathbf{J})\!+\!k_{2}\mathcal{U}^{(2)}_{k_{1}k_{2}}\!(\mathbf{J})}{k_{1}(k_{1}\!+\!k_{2})\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}^{2}}\bigg{|}^{2},$ (62)
with the coupling functions, $\mathcal{U}^{(1)}$ and $\mathcal{U}^{(2)}$, defined in Appendix B of Fouvry (2022). In Eq. (31), the sum over the resonance vectors ${(k_{1},k_{2})}$ is restricted to those such that $k_{1}$, $k_{2}$, as well ${k_{1}\!+\!k_{2}}$ are all non-zero. Importantly, we emphasise that ${|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}}$ is independent of $N$ and $q$.
Small-scale contributions. In Fig. 11, we present the contributions from small scales to the ${1/N^{2}}$ relaxation rate. Starting from Eq. (31), let us place ourselves in the limit where the three actions, $\mathbf{J}$, involved in the three-body resonance condition are close to one another, and assume that ${k_{1},k_{2}\!\simeq\!k}$. Then, from eq. (B1) in Fouvry (2022), we find that ${\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!k^{2}U_{k}^{2}}$, where we assumed that computing gradients of the form ${\partial_{J}U_{k_{1}}(J,J^{\prime})}$ does not lead to any additional scaling dependence with respect to $k$. Imposing ${U_{k}\!\simeq\!k^{-1}}$ following Eq. (45b), we obtain ${\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!1}$ as a function of $k$. Similarly, starting from eq. (B2) in Fouvry (2022), we find ${\mathcal{U}^{(2)}_{k_{1}k_{2}}\!\simeq\!k^{-1}}$. With these two scalings, we find from Eq. (62) that ${|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}\!\simeq\!k^{-4}}$. Injecting this dependence in Eq. (31), and using the definition from Eq. (35), we get ${\mathcal{R}_{2}(J,k_{1},k_{2})\!\simeq\!k^{-3}}$. Finally, we may perform the sum over ${k_{1},k_{2}}$ in Eq. (35), with the constraint ${\mathrm{Max}[|k_{1}|,|k_{2}|]\!=\!k}$, to obtain ${\mathcal{R}_{2}(J,k)\!\simeq\!k^{-2}}$. This is the scaling that is indeed recovered in Fig. 11. As expected, in that figure, we also recover that softening accelerates the decay of the contributions stemming from small scales.
Computing the kinetic prediction. In order to evaluate Eq. (31), we need to compute an integral of the form
[…]
where the integration boundaries ${J_{\pm}\!=\!J_{0}\!\pm\!\sigma_{0}}$ follow from Eq. (6). This integral can be rewritten as
[…]
It subsequently becomes
[…]
Appendix G Numerical simulations
Equations of motion. Following Eq. (8), the equations of motion of the point vortices are given by
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}=\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{ext}}+\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{self}}.$ (66)
Here, ${\mathrm{d}\mathbf{r}/\mathrm{d}t|_{\mathrm{ext}}}$ is the contribution from the background external potential, ${U_{\mathrm{ext}}}$, in Eq. (8). Since it is axisymmetric, following Eq. (5), we can compute ${\Omega_{i}\!=\!\Omega(J_{i})\!=\!\mathrm{d}U_{\mathrm{ext}}/\mathrm{d}J_{i}}$, and we find
$\frac{\mathrm{d}x_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{ext}}{}$ $=\;\;\;\Omega_{i}\,y_{i},$ (67a)
$\frac{\mathrm{d}y_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{ext}}{}$ $=-\Omega_{i}\,x_{i}.$ (67b)
For the self-interaction part, ${\mathrm{d}\mathbf{r}_{i}/\mathrm{d}t|_{\mathrm{self}}}$, following Eq. (4), we obtain
$\frac{\mathrm{d}x_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{self}}{}$ $=-\frac{G}{2\pi}\sum_{j\neq i}\gamma_{j}\,\frac{y_{i}\!-\!y_{j}}{(x_{i}\!-\!x_{j})^{2}\!+\!(y_{i}\!+\!y_{j})^{2}\!+\!\epsilon^{2}},$ (68a)
$\frac{\mathrm{d}y_{i}}{\mathrm{d}t}\bigg{|}_{\mathrm{self}}{}$ $=\;\;\;\frac{G}{2\pi}\sum_{j\neq i}\gamma_{j}\,\frac{x_{i}\!-\!x_{j}}{(x_{i}\!-\!x_{j})^{2}\!+\!(y_{i}\!-\!y_{j})^{2}\!+\!\epsilon^{2}}.$ (68b)
In practice, we fixed the softening length to ${\epsilon/J_{0}\!=\!0.01}$ in all the simulations considered. The present system has two global invariants, the total energy, ${E_{\mathrm{tot}}\!=\!H_{\mathrm{tot}}}$, as well as the total momentum, ${L_{\mathrm{tot}}\!=\!\sum_{i=1}^{N}\!\gamma_{i}\,r_{i}^{2}}$. We emphasise that evaluating Eq. (68) is costly, since it requires to perform ${\mathcal{O}(N^{2})}$ computations per evaluation of the rates of change.
Time integration. Given that the Hamiltonian from Eq. (8) is not under a straightforward separable form, we perform the time integration using a standard explicit Runge–Kutta method. Yet, since both the external potential and the pairwise interaction potentials are smooth, we may resort to high-order methods. This allows for the use of a large individual timestep making long-time integrations easier to perform. In that view, we used the 15-stage 9th-order “more efficient” Vern9 method available through the library OrdinaryDiffEq.jl.
Simulation setups. Given that we are going to perform numerical simulations over some large ranges in $N$ and $q$, it is important to carefully pick the parameters of the simulations.
[…]
For profile (2), we fixed the timestep to ${h/T_{\mathrm{dyn}}\!\simeq\!1.90\!\times\!10^{-2}}$. Each simulation was integrated up to ${T_{\mathrm{max}}/T_{\mathrm{dyn}}\!\simeq\!3.040\!\times\!10^{6}\,(0.0005/q)^{4}(N/200)^{2}}$, to match the scaling from Eq. (32). Similarly, we made 2 000 dumps of each realisation as the integration was performed. For that profile, the final relative errors on $E_{\mathrm{tot}}$ and $L_{\mathrm{tot}}$ were of order $10^{-9}$. In practice, the case ${(N,q)\!=\!(200,0.0005)}$ presented in Fig. 12 required ${\!\sim\!21\mathrm{h}}$ of computation time, per realisation, on a single core.
[…]
Figure 15: Time evolution of the number of particles, ${N(\!<\!J,t)}$, in numerical simulations, when averaged over all available realisations, for profile (1) [top] and profile (2) [bottom]. Different colors correspond to different actions $J$. The linear dependence of these time series is the sign of the onset of relaxation.
In that figure, we note that the relative change in the fraction of particles is of order $10^{-3}$, i.e. we purposely limited our measurements to the onset of relaxation. Once the average time series ${t\!\mapsto\!\langle N(\!<\!J,t)\rangle}$ determined, we performed a linear fit of the form
$\langle N(\!<\!J,t)\rangle\simeq\alpha+\beta t.$ (69)
Finally, we use the slope $\beta$ in Eqs. (17) and (33) to estimate the relaxation rates. Given the computing cost associated with each such realisation, the number of independent realisations, as well as the large number of values of ${(N,q)}$ considered, obtaining Fig. 8 was no light numerical endeavour. Finally, to obtain Fig. 9, we binned the action domain in bins of width ${10^{-2}\,J_{0}}$.
Bootstraps. In practice, it is essential to estimate the uncertainties associated with these measurements. To that purpose, we use bootstraps. A given bootstrap proceeds as follows. For every value of ${(N,q)}$, we perform the ensemble average of ${N(\!<\!J,t)}$ over a sample of $N_{\mathrm{real}}$ realisations drawn, with repetitions, from the $N_{\mathrm{real}}$ realisations available. Using the associated ${\langle N(\!<\!J,t)\rangle}$, we perform the linear fit from Eq. (69) from which we also estimate the variance of the slope $\beta$. Finally, for each ${(N,q)}$ and each considered action, we can draw a value of $\beta$ according to the associated Gaussian distribution. In practice, we performed a total of 1 000 bootstrap measurements, with which we could obtain the 16% and 84% level contours presented in Figs. 5 and 12, as well as the median values presented in Figs. 7 and 13.
Appendix H Heuristic estimate of the diffusion coefficient
When derived from the BBGKY hierarchy at order ${1/N}$, the Landau equation generically reads Chavanis (2013)
$\frac{\partial F}{\partial t}={}$ $\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\,\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{2\pi}\!\!\!\mathrm{d}\theta_{1}\,$
×$\times\big{\langle}\mathcal{F}(1\rightarrow 0,t)\,G(t,t-\tau)\,\mathcal{F}(1\rightarrow 0,t-\tau)\big{\rangle}$
×$\times\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)\,\frac{1}{\gamma}F(J_{1},t)\bigg{]},$ (70)
with
$\mathcal{F}(1\rightarrow 0,s)=-\mathrm{i}\gamma\sum_{k}k\,\mathrm{e}^{\mathrm{i}k[\theta(s)-\theta_{1}(s)]}\,U_{k}(J,J_{1}),$ (71)
the force (velocity) created by particle (vortex) $1$ on particle (vortex) 00 at time $s$, and ${G(t,t\!-\!\tau)}$ is an appropriate Green function taking into account the motion of the particles between $t$ and ${t\!-\!\tau}$ Chavanis (2013). In Eq. (70), the brackets indicate an average over the fluctuating trajectories of the particles (see below). To calculate the temporal correlation function in Eq. (70), one usually assumes that the particles follow the unperturbed mean field trajectories at leading order. This leads to the Landau equation (Eq. 14). However, this approximation breaks down close to an extremum of ${\Omega(J)}$, i.e. when particles have similar frequencies. In that case, one has to take into account fluctuations in the particles’ individual trajectories.
Following Dubin (2003), let us consider a simplified model where we approximate the motion of the particles with the effective stochastic equations
$\frac{\mathrm{d}J}{\mathrm{d}t}=0,\qquad\frac{\mathrm{d}\theta}{\mathrm{d}t}=\Omega(J)+\sqrt{2D_{\theta\theta}}\,\eta(t),$ (72)
where ${\eta(t)}$ is a Gaussian white noise such that ${\langle\eta(t)\rangle\!=\!0}$ and ${\langle\eta(t)\,\eta(t^{\prime})\rangle\!=\!\delta_{\mathrm{D}}(t\!-\!t^{\prime})}$, and $D_{\theta\theta}$ is a diffusion coefficient in angles. Integrating these equations between $t$ and ${t\!-\!\tau}$, substituting the stochastic trajectories ${J(t\!-\!\tau)\!=\!J}$ and ${\theta(t\!-\!\tau)\!=\!\theta\!-\!\Omega(J)\tau\!-\!\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}\eta(t\!-\!t^{\prime})}$ into Eq. (70), and carrying out the integration over $\theta_{1}$, we obtain
$\frac{\partial F}{\partial t}={}$ $4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
×$\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\left\langle\mathrm{e}^{\mathrm{i}k\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}[\eta(t-t^{\prime})-\eta_{1}(t-t^{\prime})]}\right\rangle$
×$\times\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$ (73)
where we introduced the frequency difference, ${\Delta\Omega(J,J_{1})\!=\!\Omega(J)\!-\!\Omega(J_{1})}$. If we assume a Gaussian distribution for the fluctuations Dubin (2003), we have
$\left\langle\mathrm{e}^{\mathrm{i}k\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}\eta(t-t^{\prime})}\right\rangle=\mathrm{e}^{-k^{2}D_{\theta\theta}\tau},$ (74)
and the kinetic equation (Eq. 70) becomes
$\frac{\partial F}{\partial t}{}$ $=4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (75)
×$\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\mathrm{e}^{-\tau/T_{k}}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$
where we introduced the (scale-dependent) decorrelation time, ${T_{k}\!=\!1/(2D_{\theta\theta}k^{2})}$. Integrating over $\tau$, we obtain
$\frac{\partial F}{\partial t}=4\pi^{2}\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (76)
×$\times\delta_{T_{k}}\big{(}k\,\Delta\Omega[J,J_{1}]\big{)}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$
with ${\delta_{T_{k}}(\omega)}$ the broadened Dirac delta from Eq. (24). [^10] In the main text, we assumed ${T_{k}\!=\!T_{\mathrm{reg}}/k}$. This leads to the identity ${\delta_{T_{k}}(k\,\omega)\!=\!\delta_{T_{\mathrm{reg}}}(\omega)/|k|}$, similar to the one satisfied by the usual Dirac delta. From this identity results the broadened Landau equation (Eq. 25). Here, the function ${\delta_{T_{k}}(k\,\omega)}$ has a different dependence with $k$, so the broadened Landau equation (Eq. 76) is slightly different. See Chavanis and Lemou (2007) for another regularization of the Dirac delta. Explicitly, the kinetic equation reads
$\frac{\partial F}{\partial t}=4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$ (77)
×$\times\frac{T_{k}}{1\!+\!(k\,\Delta\Omega[J,J_{1}]\,T_{k})^{2}}\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]}.$
If we let ${T_{k}\!\rightarrow\!+\infty}$, or if ${k\,\Delta\Omega\,T_{k}\!\gg\!1}$, which both amount to neglecting the fluctuations in particles individual trajectories, Eq. (77) becomes the Landau equation (Eq. 14).
However, when $J$ is close to the extremum of ${\Omega(J)}$, which we abusively take to be in $J_{0}$, we are faced with the opposite limit, namely ${k\,\Delta\Omega\,T_{k}\!\ll\!1}$. In that case, the kinetic equation (Eq. 77) reduces to
$\frac{\partial F}{\partial t}={}$ $4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
×$\times T_{k}\,\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]}.$ (78)
The (broadened) Landau equation (Eq. 76) can be viewed as a Fokker–Planck equation in action space involving a diffusion term and a drift term. The diffusion coefficient in action reads
$D=4\pi^{2}\gamma\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
×$\times\,\delta_{T_{k}}\big{(}k\,\Delta\Omega[J,J_{1}]\big{)}F(J_{1},t).$ (79)
As expected, this equation reduces to the usual Landau diffusion coefficient when ${T_{k}\!\rightarrow\!+\infty}$, and scales like ${1/N}$ (via the prefactor $\gamma$). However, if we assume ${k\,\Delta\Omega\,T_{k}\!\ll\!1}$ and heuristically identify the diffusion coefficient in angles, $D_{\theta\theta}$, with the diffusion coefficient from Eq. (79), i.e. if we write ${D_{\theta\theta}\!\simeq\!D/J_{0}^{2}}$, we obtain for ${J\!\rightarrow\!J_{0}}$
$D^{2}=2\pi\gamma\,J_{0}^{2}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\sum_{k>0}|U_{k}(J,J_{1})|^{2}F(J_{1},t).$ (80)
Since ${\gamma\!\sim\!1/N}$, we recover the Taylor–McNamara scaling Taylor and McNamara (1971), ${D\!\sim\!1/\sqrt{N}}$. If we account for an active fraction of particles, as in the main text, this scaling becomes ${D\!\sim\!q/\sqrt{N}}$. We note that the diffusion coefficient from Eq. (80) could be obtained numerically, and subsequently compared with measurements in numerical simulations.
Along the same line, if we make the local approximation, ${F(J_{1},t)\!\simeq\!F(J_{0},t)}$, the integral in Eq. (80) can be estimated analytically when using the (unsoftened) interaction potential from Eq. (43). Overall, Eq. (80) leads to the rough estimate
$D^{2}\sim\frac{J_{0}^{3}}{8\pi}\,\gamma\,F(J_{0},t),$ (81)
up to logarithmic corrections. To make a closer connection with the results of Taylor–McNamara Taylor and McNamara (1971), we introduce the original diffusion coefficient in radius $D_{rr}$, related to $D$ by ${D\!=\!D_{rr}r^{2}\!=\!2JD_{rr}}$, as well as the typical total circulation ${\Gamma\!\sim\!2\pi F(J_{0},t)J_{0}}$. We obtain
$D_{rr}^{2}\sim\frac{J_{0}}{32\pi}\,\gamma\,F(J_{0},t)\sim\frac{1}{64\pi^{2}}\,\Gamma\gamma\,\sim\frac{1}{64\pi^{2}}\,N\,\gamma^{2}.$ (82)
Of course, following our numerous heuristic approximations, the prefactor of Eq. (82) is highly uncertain. Nonetheless, this should be compared with the Taylor–McNamara formula ${D_{rr}^{2}\!=\!N\gamma^{2}/(64\pi^{3})}$ Taylor and McNamara (1971). More generally, Eq. (79) yields a self-consistent equation for the diffusion coefficient which connects the Landau form (${k\Delta\Omega T_{k}\!\gg\!1}$) and the Taylor–McNamara form (${k\Delta\Omega T_{k}\!\ll\!1}$).